# FDS: Feedback-guided Domain Synthesis with Multi-Source Conditional Diffusion Models for Domain Generalization

### v2

@@ 6.2 Further Analysis
Table 2: Comparative analysis of FDS component effects on accuracy (%) across PACS dataset domains. "Basic" refers to generation without interpolation or filtering.
Module Target Domains Art Cartoon Photo Sketch Avg. Baseline (SWAD [24]) 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$1.0 88.11 $\pm$0.4 + Basic 89.87 $\pm$0.1 85.59 $\pm$0.6 97.50 $\pm$0.3 83.07 $\pm$0.4 89.01 $\pm$0.4 + Interpolation 91.38 $\pm$0.2 85.20 $\pm$0.6 97.73 $\pm$0.1 84.27 $\pm$0.9 89.65 $\pm$0.4 + Filtering 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50 $\pm$0.3
Ablation Study on Different Components.  To evaluate the impact of each component on generalization, we systematically introduce each module and observe the enhancements. As detailed in Table 2, first incorporating domain-specific synthetic samples, generated without interpolation or filtering (Basic), yields a $0.9\%$ gain over the baseline. This increment validates our diffusion model’s proficiency in capturing and replicating the class-specific distributions within domains, thus refining OOD performance through enhanced density estimation near original samples. Subsequently, integrating images from pseudo-novel domains via our interpolation strategy leads to a further $0.5\%$ enhancement, underscoring the mechanism’s effectiveness in connecting distinct domains. Lastly, applying our entropy-based filtering to eliminate overly simplistic images and images with diminished semantic relevance results in a significant $2.4\%$ improvement against the SWAD model, a robust benchmark. These outcomes collectively underscore the efficacy of our approach in improving OOD generalization.
@@
Figure 5: t-SNE plots showcasing the original and generated samples for the "Art", "Photo", and "Sketch" source domains and the "giraffe" class in the PACS dataset using our method. Features were extracted using the CLIP vision encoder [71].
@@
Filtering Mechanism.  One might consider that the enhanced accuracy observed with our filtering mechanism might not solely be its merit but rather the result of constraining the sample size for a balanced final training set. To investigate this hypothesis, Figure 4 contrasts the outcome of selecting $N_{L}$ samples at random from all generated images per class against employing our entropy-based filtering strategy. The consistent improvement in out-of-domain (OOD) generalization across all domains, facilitated by our filtering approach, underscores its effectiveness.
@@
Table 5: Domain diversity metric [72] between source domains and the target domain of the PACS dataset. "Basic" refers to generation without interpolation or filtering.
@@
Data Target Domains Art Cartoon Photo Sketch Avg. Original PACS 0.39 0.87 0.54 0.99 0.70 Basic 0.54 0.72 0.45 1.01 0.68 FDS 0.44 0.58 0.49 0.92 0.61
@@
Domain Diversity Visualization.  To illustrate the effectiveness of our method, we present t-SNE plots of original PACS dataset samples and those generated by FDS in Figure 5. The t-SNE plots show distinct clusters for original source domains (here Art, Photo, Sketch) and highlight how the generated samples clearly bridge the gaps between these clusters, thus enhancing domain diversity. This expanded diversity is crucial for improving the generalization capabilities of models, as it ensures a broader spectrum of data distributions in the training set. By providing a continuous representation of the domain space, our method facilitates smoother transitions and better prepares models to handle unseen domains, ultimately contributing to more robust performance in real-world applications.
Domain Diversity Quantification.  We further validate the effectiveness of our method using a domain diversity metric based on the methodology proposed in [72]. This metric quantifies the diversity shift between the source domains and the target domain, providing valuable insight into how the newly generated domains using our method can improve generalization on unseen domains. Table 5 presents the domain diversity metric for the original PACS dataset, comparing it with samples generated using the diffusion model in its basic form (no interpolation and no filter) and our FDS method (including interpolation and filtering). This table demonstrates that our FDS method results in a lower diversity metric compared to using the original PACS dataset or the samples generated with basic generation, suggesting that our method more effectively promotes generalization to new, unseen domains.
@@
Table 6: Impact of FDS on in-domain PACS accuracy (%). ’A’, ’C’, ’P’, and ’S’ show ’Art’, ’Cartoon’, ’Photo’, and ’Sketch’ domains.
@@
Method Source Domains C, S, P A, P, S A, C, S A, C, P Avg. ERM [44] 97.72 $\pm$0.2 97.72 $\pm$0.2 96.54 $\pm$0.2 97.46 $\pm$0.4 97.36 $\pm$0.2 ERM + FDS (ours) 98.14 $\pm$0.4 97.86 $\pm$0.5 97.44 $\pm$0.2 98.04 $\pm$0.1 97.87 $\pm$0.2 SWAD [24] 98.23 $\pm$0.2 97.95 $\pm$0.4 97.30 $\pm$0.1 98.40 $\pm$0.2 97.97 $\pm$0.2 SWAD + FDS (ours) 98.43 $\pm$0.4 98.12 $\pm$0.1 97.90 $\pm$0.3 98.72 $\pm$0.2 98.29 $\pm$0.1
In-Domain Regularization Effect.  In this section, we study the impact of incorporating generated images produced by our method on the accuracy of the network when tested on the same domains on which it was trained. As indicated in Table 6, our FDS approach surpasses the baseline accuracy within this context. This affirms the value of our method, not only in OOD conditions, but also in standard in-domain validation, aligning with the principles of Vicinal Risk Minimization (VRM) [57]. Hence, our method may also be viewed as a regularization strategy, suitable for a broad spectrum of applications.
@@ 7 Conclusion
This work presented FDS, a domain generalization (DG) technique that leverages diffusion models for domain mixing, generating a diverse set of images to bridge the domain gap between source domains distribution. We also proposed an entropy-based filtering strategy enriches the pseudo-novel generated set with images that test the limits of classifiers trained on original data, thereby boosting generalization. Our extensive experiments across multiple benchmarks demonstrate that our method not only surpasses existing DG techniques but also sets new records for accuracy. Our analysis indicate that our approach contributes to more stable training processes when confronted with domain shifts and serves effectively as a regularization method in in-domain contexts. Notably, our technique consistently enhances performance across diverse scenarios, from realistic photos to sketches. Moreover, while we exploited our trained diffusion model for covering the domain gap, more sophisticated techniques could be considered. For instance, future work could investigate the idea of generating images by extrapolate in domain space to explore pseudo-novel distributions.
FDS: Feedback-guided Domain Synthesis with Multi-Source Conditional Diffusion Models for Domain Generalization - Supplementary Material
@@ Appendix A Implementation
Our proposed FDS method is built using the Python language and the PyTorch framework. We utlized four NVIDIA A100 GPUs for all our experiments. For initializing our models, we utilize the original Stable Diffusion version 1.5 as our initial weight  [34]. The key hyperparameter configurations employed for training these diffusion models and generating new domains are detailed in Tables 8 and 8, respectively.
Furthermore, for classifier training, we adhere to the methodologies and parameter settings described by Cha et al. [24], ensuring consistency and reproducibility in our experimental setup. The original implementation and instructions for reproducing our results are accessible via our GitHub repository.
@@ Appendix B Additional Ablation
Method Augmented Domains Accuracy (%) Art Cartoon Photo Sketch Avg. SWAD (reproduced) 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$1.0 88.11 $\pm$0.45 SWAD + FPS ID0 91.03 $\pm$0.5 83.87 $\pm$0.6 97.75 $\pm$0.3 85.77 $\pm$0.4 89.61 $\pm$0.30 SWAD + FPS ID1 91.01 $\pm$0.6 85.06 $\pm$1.3 97.90 $\pm$0.3 83.64 $\pm$0.4 89.40 $\pm$0.65 SWAD + FPS ID2 91.46 $\pm$0.3 85.22 $\pm$0.8 97.88 $\pm$0.2 84.27 $\pm$0.3 89.71 $\pm$0.40 SWAD + FPS ID0 + ID1 91.52 $\pm$0.0 85.87 $\pm$0.7 98.03 $\pm$0.3 85.70 $\pm$1.0 90.28 $\pm$0.50 SWAD + FPS ID1 + ID2 91.62 $\pm$0.8 85.57 $\pm$0.4 98.20 $\pm$0.3 83.88 $\pm$0.6 89.82 $\pm$0.53 SWAD + FPS ID0 + ID2 91.52 $\pm$0.1 84.54 $\pm$0.5 98.28 $\pm$0.1 86.45 $\pm$0.8 90.20 $\pm$0.38 SWAD + FPS ID0 + ID1 + ID2 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50 $\pm$0.35
@@ Appendix D Detailed Results
Here we present the comprehensive tables containing all the detailed information that was summarized in the main paper. The leave-one-out performance (train-domain validation set) across different domains of PACS, VLCS, and OfficeHome datasets are detailed in the tables 12, 13, and 14, respectively. Additionally, the oracle (test-domain validation set) accuracy results for the PACS, VLCS, and OfficeHome benchmarks are detailed in Table 15, 16, and 17, respectively.
@@
Method Aug. Target Domains Art Cartoon Photo Sketch Avg. Standard Methods ERM (baseline) [44] 84.7 $\pm$0.4 80.8 $\pm$0.6 97.2 $\pm$0.3 79.3 $\pm$1.0 85.5 $\pm$0.2 ERM (reproduced) 86.9 $\pm$0.6 80.2 $\pm$0.7 96.6 $\pm$0.4 74.5 $\pm$2.9 84.3 $\pm$1.1 IRM [60] 84.8 $\pm$1.3 76.4 $\pm$1.1 96.7 $\pm$0.6 76.1 $\pm$1.0 83.5 $\pm$0.8 GroupDRO [47] 83.5 $\pm$0.9 79.1 $\pm$0.6 96.7 $\pm$0.3 78.3 $\pm$2.0 84.4 $\pm$0.8 Mixup [61] 86.1 $\pm$0.5 78.9 $\pm$0.8 97.6 $\pm$0.1 75.8 $\pm$1.8 84.6 $\pm$0.6 CORAL [62] 88.3 $\pm$0.2 80.0 $\pm$0.5 97.5 $\pm$0.3 78.8 $\pm$1.3 86.2 $\pm$0.3 MMD [63] 86.1 $\pm$1.4 79.4 $\pm$0.9 96.6 $\pm$0.2 76.5 $\pm$0.5 84.6 $\pm$0.5 DANN [64] 86.4 $\pm$0.8 77.4 $\pm$0.8 97.3 $\pm$0.4 73.5 $\pm$2.3 83.6 $\pm$0.4 MLDG [15] 85.5 $\pm$1.4 80.1 $\pm$1.7 97.4 $\pm$0.3 76.6 $\pm$1.1 84.9 $\pm$1.1 VREx [74] 86.0 $\pm$1.6 79.1 $\pm$0.6 96.9 $\pm$0.5 77.7 $\pm$1.7 84.9 $\pm$1.1 ARM [75] 86.8 $\pm$0.6 76.8 $\pm$0.5 97.4 $\pm$0.3 79.3 $\pm$1.2 85.1 $\pm$0.6 SagNet [65] 87.4 $\pm$1.0 80.7 $\pm$0.6 97.1 $\pm$0.1 80.0 $\pm$0.4 86.3 $\pm$0.2 RSC [23] 85.4 $\pm$0.8 79.7 $\pm$1.8 97.6 $\pm$0.3 78.2 $\pm$1.2 85.2 $\pm$0.9 Mixstyle [20] 86.8 $\pm$0.5 79.0 $\pm$1.4 96.6 $\pm$0.1 78.5 $\pm$2.3 85.2 $\pm$0.3 mDSDI [66] 87.7 $\pm$0.4 80.4 $\pm$0.7 98.1 $\pm$0.3 78.4 $\pm$1.2 86.2 $\pm$0.2 SelfReg [41] 87.9 $\pm$1.0 79.4 $\pm$1.4 96.8 $\pm$0.7 78.3 $\pm$1.2 85.6 $\pm$0.4 Fishr [73] 88.4 $\pm$0.2 78.7 $\pm$0.7 97.0 $\pm$0.1 77.8 $\pm$2.0 85.5 $\pm$0.5 DCAug [59] 88.5 $\pm$0.8 78.8 $\pm$1.5 96.3 $\pm$0.1 80.8 $\pm$0.5 86.1 $\pm$0.7 CDGA [55] 89.1 $\pm$1.0 82.5 $\pm$0.5 97.4 $\pm$0.2 84.8 $\pm$0.9 88.5 $\pm$0.5 ERM + FDS (ours) 90.7 $\pm$0.9 84.2 $\pm$0.6 97.2 $\pm$0.1 83.0 $\pm$0.4 88.8 $\pm$0.1 WA Methods SWAD (baseline) [24] 89.3 $\pm$0.2 83.4 $\pm$0.6 97.3 $\pm$0.3 82.5 $\pm$0.5 88.1 $\pm$0.1 SWAD (reproduced) 89.5 $\pm$0.2 83.7 $\pm$0.4 97.3 $\pm$0.2 82.1 $\pm$0.1 88.1 $\pm$0.4 SelfReg SWA [41] 85.9 $\pm$0.6 81.9 $\pm$0.4 96.8 $\pm$0.1 81.4 $\pm$0.6 86.5 $\pm$0.3 DNA [67] 89.8 $\pm$0.2 83.4 $\pm$0.4 97.7 $\pm$0.1 82.6 $\pm$0.2 88.4 $\pm$0.1 DiWA [68] 90.1 $\pm$0.6 83.3 $\pm$0.6 98.2 $\pm$0.1 83.4 $\pm$0.4 88.8 $\pm$0.4 TeachDCAug [59] 89.6 $\pm$0.0 81.8 $\pm$0.5 97.7 $\pm$0.0 84.5 $\pm$0.2 88.4 $\pm$0.2 SWAD + FDS (ours) 91.8 $\pm$0.3 86.0 $\pm$0.8 98.1 $\pm$0.2 86.1 $\pm$0.1 90.5 $\pm$0.3

### v3

@@ 6.2 Further Analysis
Ablation Study on Different Components.  To evaluate the impact of each component on generalization, we systematically introduce each module and observe the enhancements. As detailed in Table 2, first incorporating domain-specific synthetic samples, generated without interpolation or filtering (Basic Gen.), yields a $0.9\%$ gain over the baseline. This increment validates our diffusion model’s proficiency in capturing and replicating the class-specific distributions within domains, thus refining OOD performance through enhanced density estimation near original samples. Subsequently, integrating images from pseudo-novel domains via our interpolation strategy leads to a further $0.5\%$ enhancement, underscoring the mechanism’s effectiveness in connecting distinct domains. Lastly, applying our entropy-based filtering to eliminate overly simplistic images and images with diminished semantic relevance results in a significant $2.4\%$ improvement against the SWAD model, a robust benchmark. These outcomes collectively underscore the efficacy of our approach in improving OOD generalization.
@@
Figure 5: t-SNE plots showcasing the original “giraffe” class samples for the “Art”, “Photo” and “Sketch” source domains of the PACS dataset, as well as the data generated with FDS.
@@
Filtering Mechanism.  One might consider that the enhanced accuracy observed with our filtering mechanism might be the result of constraining the sample size for a balanced final training set. To investigate this hypothesis, Figure 4 contrasts the outcome of selecting $N_{L}$ samples at random from all generated images per class against employing our entropy-based filtering strategy. The consistent improvement in out-of-domain (OOD) generalization across all domains, facilitated by our filtering approach, underscores its effectiveness.
@@
Data Target Domains Art Cartoon Photo Sketch Avg. Original PACS 0.39 0.87 0.54 0.99 0.70 Basic Gen. 0.54 0.72 0.45 1.01 0.68 FDS 0.44 0.58 0.49 0.92 0.61
Table 5: Domain diversity metric [72] between source domains and the target domain of the PACS dataset. “Basic Gen.” refers to generation without interpolation or filtering.
@@
Domain Diversity Visualization.  To illustrate the effectiveness of our method, we present t-SNE plots of original PACS dataset samples and those generated by FDS in Figure 5. The t-SNE plots show distinct clusters for original source domains (here Art, Photo, Sketch) and highlight how the generated samples clearly bridge the gaps between these clusters, thus enhancing domain diversity. This expanded diversity is crucial for improving the generalization capabilities of models, as it ensures a broader spectrum of data distributions in the training set. By providing a continuous representation of the domain space, our method facilitates smoother transitions and better prepares models to handle unseen domains, ultimately contributing to more robust performance in real-world applications. Please check the supplementary materials for further details.
Method Source Domains C, S, P A, P, S A, C, S A, C, P Avg. ERM 98.05 $\pm$0.5 96.06 $\pm$0.9 95.66 $\pm$0.3 96.83 $\pm$0.6 96.65 ERM + Dup. Aug. 98.01 $\pm$0.2 97.07 $\pm$0.3 96.46 $\pm$0.2 97.28 $\pm$0.3 97.21 ERM + Basic Gen. 96.96 $\pm$0.4 96.74 $\pm$0.2 96.23 $\pm$0.2 97.37 $\pm$0.5 96.83 ERM + FDS (ours) 97.77 $\pm$0.4 97.53 $\pm$0.5 96.71 $\pm$0.1 98.02 $\pm$0.4 97.51 SWAD 98.48 $\pm$0.3 97.87 $\pm$0.5 97.61 $\pm$0.2 98.44 $\pm$0.0 98.10 SWAD + Dup. Aug. 98.35 $\pm$0.2 97.86 $\pm$0.3 97.38 $\pm$0.1 98.16 $\pm$0.1 97.94 SWAD + Basic Gen. 98.42 $\pm$0.4 98.23 $\pm$0.3 97.87 $\pm$0.5 98.64 $\pm$0.2 98.29 SWAD + FDS (ours) 98.70 $\pm$0.3 98.25 $\pm$0.5 97.80 $\pm$0.4 98.61 $\pm$0.0 98.34
@@
Table 6: Impact of FDS on in-domain PACS accuracy (%). ‘A’, ‘C’, ‘P’, and ‘S’ refer to ‘Art’, ‘Cartoon’, ‘Photo’, and ‘Sketch’.
@@
Method Target Domains Art Cartoon Photo Sketch Avg. ERM 86.94 $\pm$0.6 80.21 $\pm$0.7 96.61 $\pm$0.4 74.45 $\pm$2.9 84.30 ERM + Dup. Aug. 85.34 $\pm$0.9 80.99 $\pm$0.9 94.98 $\pm$0.8 76.88 $\pm$2.0 84.55 ERM + Basic Gen. 87.21 $\pm$0.3 80.90 $\pm$1.9 95.71 $\pm$0.4 80.31 $\pm$2.3 86.03 ERM + FDS (ours) 90.69 $\pm$0.9 84.19 $\pm$0.6 97.21 $\pm$0.1 82.99 $\pm$0.4 88.77 SWAD 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$1.0 88.11 SWAD + Dup. Aug. 89.55 $\pm$0.3 83.00 $\pm$1.7 97.65 $\pm$0.2 81.86 $\pm$1.1 88.02 SWAD + Basic Gen. 89.87 $\pm$0.1 85.59 $\pm$0.6 97.50 $\pm$0.3 83.07 $\pm$0.4 89.01 SWAD + FDS (ours) 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50
Table 7: Leave-one-out accuracy (%) of FDS compared to other augmentation strategies.
Domain Diversity Quantification.  We further validate the effectiveness of our method using a domain diversity metric based on the methodology proposed in [72]. This metric quantifies the diversity shift between the source domains and the target domain, providing valuable insight into how the newly generated domains using our method can improve generalization on unseen domains. Table 5 presents the domain diversity metric for the original PACS dataset, comparing it with samples generated using the diffusion model in its basic form (no interpolation and no filter) and our FDS method (including interpolation and filtering). We observe that our FDS method results in a lower diversity metric compared to using the original PACS dataset or the samples generated with basic generation, suggesting that our method more effectively promotes generalization to new, unseen domains.
In-Domain Regularization Effect.  In this section, we study the impact of incorporating images generated by our method on the network’s accuracy, when testing on the same domains as those used for training. To provide a more comprehensive analysis, we adopted an 80/10/10 split for the source domains222 This setting was only used for in-domain experiments, while the rest of the paper follows the standard 80/20 setting as used in DomainBed.. We selected the best model on the validation set and reported its performance on the held-out in-domain test set. As indicated in Table 6, our FDS approach surpasses the baseline accuracy in this setup for both the ERM and SWAD backbones. Additionally, we compared our method to two alternative strategies: i) duplicating and augmenting original samples (Dup. Aug.), and ii) generating synthetic in-domain images (same number of images as FDS) without domain interpolation (Basic Gen.). While these strategies also improve the performance, FDS achieves the highest accuracy due to its ability to generate more diverse and challenging pseudo-domains, rather than simply increasing dataset size. This confirms the value of our method not only in OOD conditions but also in standard in-domain validation, aligning with the principles of Vicinal Risk Minimization (VRM) [58]. Hence, our method may also be viewed as a regularization strategy, suitable for a broad spectrum of applications. For completeness, we also report the out-domain performance of FDS and these strategies in Table 7. As can be seen, the beneficial impact of FDS on OOD generalization is not simply the result of increasing the dataset size.
@@ 7 Conclusion
This work presented FDS, a domain generalization (DG) technique that leverages diffusion models for domain mixing, generating a diverse set of images to bridge the domain gap between source domains distribution. We also proposed an entropy-based filtering strategy that enriches the pseudo-novel generated set with images that test the limits of classifiers trained on original data, thereby boosting generalization. Our extensive experiments across multiple benchmarks demonstrate that our method not only surpasses existing DG techniques but also sets new records for accuracy. Our analysis indicate that our approach contributes to more stable training processes when confronted with domain shifts and serves effectively as a regularization method in in-domain contexts. Notably, our technique consistently enhances performance across diverse scenarios, from realistic photos to sketches. While we exploited our trained diffusion model for covering the domain gap, more sophisticated techniques could be considered. For instance, future work could investigate the idea of generating pseudo-novel distributions via extrapolation in the domain space, in addition to interpolation.
FDS: Feedback-guided Domain Synthesis with Multi-Source Conditional Diffusion Models for Domain Generalization – Supplementary Material
@@ Appendix A Implementation
Our proposed FDS method is built using the Python language and the PyTorch framework. We utlized four NVIDIA A100 GPUs for all our experiments. For initializing our models, we utilize the original Stable Diffusion version 1.5 as our initial weight  [34]. The key hyperparameter configurations employed for training these diffusion models and generating new domains are detailed in Tables 8 and 9, respectively.
Furthermore, for classifier training, we adhere to the methodologies and parameter settings described by Cha et al. [23], ensuring consistency and reproducibility in our experimental setup. The original implementation and instructions for reproducing our results are accessible via https://github.com/Mehrdad-Noori/FDS.
@@ Appendix B Additional Ablation
Method Augmented Domains Accuracy (%) Art Cartoon Photo Sketch Avg. SWAD (reproduced) 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$1.0 88.11 $\pm$0.45 SWAD + FDS ID0 91.03 $\pm$0.5 83.87 $\pm$0.6 97.75 $\pm$0.3 85.77 $\pm$0.4 89.61 $\pm$0.30 SWAD + FDS ID1 91.01 $\pm$0.6 85.06 $\pm$1.3 97.90 $\pm$0.3 83.64 $\pm$0.4 89.40 $\pm$0.65 SWAD + FDS ID2 91.46 $\pm$0.3 85.22 $\pm$0.8 97.88 $\pm$0.2 84.27 $\pm$0.3 89.71 $\pm$0.40 SWAD + FDS ID0 + ID1 91.52 $\pm$0.0 85.87 $\pm$0.7 98.03 $\pm$0.3 85.70 $\pm$1.0 90.28 $\pm$0.50 SWAD + FDS ID1 + ID2 91.62 $\pm$0.8 85.57 $\pm$0.4 98.20 $\pm$0.3 83.88 $\pm$0.6 89.82 $\pm$0.53 SWAD + FDS ID0 + ID2 91.52 $\pm$0.1 84.54 $\pm$0.5 98.28 $\pm$0.1 86.45 $\pm$0.8 90.20 $\pm$0.38 SWAD + FDS ID0 + ID1 + ID2 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50 $\pm$0.35
@@
t-SNE Visualizations.  This section presents a comprehensive t-SNE analysis for all classes in the PACS dataset, demonstrating the effectiveness of the FDS method in generating diverse, high-quality samples. The plots are provided in Figure 13. Each t-SNE plot illustrates the distribution of both original and FDS-generated samples across different domains. The results shown here are based on our diffusion model trained on the “Art,” “Photo,” and “Sketch” source domains from the PACS dataset. To create these visualizations, we extracted features using the CLIP vision encoder [73]. Each class in the PACS dataset is represented as distinct clusters, with “x” markers indicating the location of the average representation of each domains. These averages serve as a reference to assess how well the FDS-generated samples are compared with the original domains. These plots demonstrate how FDS enables smooth transitions between domains by interpolating between domain characteristics. This ability to generate synthetic data across a broad spectrum of domain representations improves the diversity of training data and enhances the model’s generalization ability. By covering a wider range of the domain space, FDS helps the model better handle unseen domains, making it more robust in real-world applications. These visualizations also suggest that the generated domains can be viewed as new pseudo-domains, as the FDS samples exhibit distributions distinct from their original sources domains. This additional diversity is critical for training models capable of generalizing beyond the source domains.
Visual Comparisons.  This section visually compares the original images from the PACS dataset with the synthetic images generated by our FDS method, highlighting the ability of FDS to interpolate between domains. We provide examples for each pair of source domains used in training: "Art," "Photo," and "Sketch.". The visual comparisions are illustrated in Figures 14, 15, and 16. Each figure contains three sections: the first section shows samples from one original PACS domain, the middle section contains FDS-generated images combining the two selected domains, and the final section shows samples from the other original domain. These visual comparisons show that the FDS-generated images effectively blend domain-specific features, offering new pseudo-domain that can enrich the training set and enhance model generalization.
@@ Appendix D Detailed Results
Here we present the comprehensive tables containing all the detailed information that was summarized in the main paper. The leave-one-out performance (train-domain validation set) across different domains of PACS, VLCS, and OfficeHome datasets are detailed in the tables 13, 14, and 15, respectively. Additionally, the oracle (test-domain validation set) accuracy results for the PACS, VLCS, and OfficeHome benchmarks are detailed in Table 16, 17, and 18, respectively.
Method Aug. Target Domains Art Cartoon Photo Sketch Avg. Standard Methods ERM (baseline) [44] 84.7 $\pm$0.4 80.8 $\pm$0.6 97.2 $\pm$0.3 79.3 $\pm$1.0 85.5 $\pm$0.2 ERM (reproduced) 86.9 $\pm$0.6 80.2 $\pm$0.7 96.6 $\pm$0.4 74.5 $\pm$2.9 84.3 $\pm$1.1 IRM [61] 84.8 $\pm$1.3 76.4 $\pm$1.1 96.7 $\pm$0.6 76.1 $\pm$1.0 83.5 $\pm$0.8 GroupDRO [47] 83.5 $\pm$0.9 79.1 $\pm$0.6 96.7 $\pm$0.3 78.3 $\pm$2.0 84.4 $\pm$0.8 Mixup [62] 86.1 $\pm$0.5 78.9 $\pm$0.8 97.6 $\pm$0.1 75.8 $\pm$1.8 84.6 $\pm$0.6 CORAL [63] 88.3 $\pm$0.2 80.0 $\pm$0.5 97.5 $\pm$0.3 78.8 $\pm$1.3 86.2 $\pm$0.3 MMD [64] 86.1 $\pm$1.4 79.4 $\pm$0.9 96.6 $\pm$0.2 76.5 $\pm$0.5 84.6 $\pm$0.5 DANN [65] 86.4 $\pm$0.8 77.4 $\pm$0.8 97.3 $\pm$0.4 73.5 $\pm$2.3 83.6 $\pm$0.4 MLDG [14] 85.5 $\pm$1.4 80.1 $\pm$1.7 97.4 $\pm$0.3 76.6 $\pm$1.1 84.9 $\pm$1.1 VREx [75] 86.0 $\pm$1.6 79.1 $\pm$0.6 96.9 $\pm$0.5 77.7 $\pm$1.7 84.9 $\pm$1.1 ARM [76] 86.8 $\pm$0.6 76.8 $\pm$0.5 97.4 $\pm$0.3 79.3 $\pm$1.2 85.1 $\pm$0.6 SagNet [66] 87.4 $\pm$1.0 80.7 $\pm$0.6 97.1 $\pm$0.1 80.0 $\pm$0.4 86.3 $\pm$0.2 RSC [22] 85.4 $\pm$0.8 79.7 $\pm$1.8 97.6 $\pm$0.3 78.2 $\pm$1.2 85.2 $\pm$0.9 Mixstyle [19] 86.8 $\pm$0.5 79.0 $\pm$1.4 96.6 $\pm$0.1 78.5 $\pm$2.3 85.2 $\pm$0.3 mDSDI [67] 87.7 $\pm$0.4 80.4 $\pm$0.7 98.1 $\pm$0.3 78.4 $\pm$1.2 86.2 $\pm$0.2 SelfReg [41] 87.9 $\pm$1.0 79.4 $\pm$1.4 96.8 $\pm$0.7 78.3 $\pm$1.2 85.6 $\pm$0.4 Fishr [74] 88.4 $\pm$0.2 78.7 $\pm$0.7 97.0 $\pm$0.1 77.8 $\pm$2.0 85.5 $\pm$0.5 DCAug [60] 88.5 $\pm$0.8 78.8 $\pm$1.5 96.3 $\pm$0.1 80.8 $\pm$0.5 86.1 $\pm$0.7 DomainDiff [54] 84.9 $\pm$1.6 82.9 $\pm$0.0 95.5 $\pm$0.0 79.0 $\pm$0.9 85.6 $\pm$0.6 DSI [55] 84.6 $\pm$2.4 81.4 $\pm$1.6 96.8 $\pm$0.5 82.5 $\pm$1.0 86.9 $\pm$1.4 CDGA [56] 89.1 $\pm$1.0 82.5 $\pm$0.5 97.4 $\pm$0.2 84.8 $\pm$0.9 88.5 $\pm$0.5 ERM + FDS (ours) 90.7 $\pm$0.9 84.2 $\pm$0.6 97.2 $\pm$0.1 83.0 $\pm$0.4 88.8 $\pm$0.1 WA Methods SWAD (baseline) [23] 89.3 $\pm$0.2 83.4 $\pm$0.6 97.3 $\pm$0.3 82.5 $\pm$0.5 88.1 $\pm$0.1 SWAD (reproduced) 89.5 $\pm$0.2 83.7 $\pm$0.4 97.3 $\pm$0.2 82.1 $\pm$0.1 88.1 $\pm$0.4 SelfReg SWA [41] 85.9 $\pm$0.6 81.9 $\pm$0.4 96.8 $\pm$0.1 81.4 $\pm$0.6 86.5 $\pm$0.3 DNA [68] 89.8 $\pm$0.2 83.4 $\pm$0.4 97.7 $\pm$0.1 82.6 $\pm$0.2 88.4 $\pm$0.1 DiWA [69] 90.1 $\pm$0.6 83.3 $\pm$0.6 98.2 $\pm$0.1 83.4 $\pm$0.4 88.8 $\pm$0.4 TeachDCAug [60] 89.6 $\pm$0.0 81.8 $\pm$0.5 97.7 $\pm$0.0 84.5 $\pm$0.2 88.4 $\pm$0.2 SWAD + FDS (ours) 91.8 $\pm$0.3 86.0 $\pm$0.8 98.1 $\pm$0.2 86.1 $\pm$0.1 90.5 $\pm$0.3